# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ 1 Introduction
Multimodal fusion is a critical and extensively studied problem in many significant domains (shaik2023survey, ; yang2007online, ; tsai2019multimodal, ; cao2023multi, ), such as sentiment analysis han2021improving ; majumder2018multimodal , image and video captioning karpathy2015deep ; johnson2016densecap , and medical prediction huang2020multimodal ; soenksen2022integrated . Previous research has shown that embracing multimodality can improve predictive performance by capturing complementary information across modalities, outperforming single-modality approaches in similar tasks (potamianos2003recent, ; huang2020fusion, ). However, an ongoing challenge lies in the creation of scalable frameworks for fusing multimodal data under a variety of conditions, and in creating reliable models that consistently surpass their single-modal counterparts.
Handling a variable number of input modalities remains an open challenge in multimodal fusion, due to challenges with scalability and lack of unified approaches for addressing missing modalities. Many existing multimodal fusion methods are designed for only two modalities han2021bi ; zhou2019review ; zhang2023improving , rely on costly pairwise comparisons between modalities tsai2019multimodal , or employ simple concatenation approaches soenksen2022integrated , rendering them unable to scale to settings with a large number of input modalities or adequately capture inter-modal interactions. Similarly, existing works are either unable to handle missing modalities entirely zhang2023improving ; zhan2021product1m or use imputation approaches tran2017missing ; liu2023attention ; soenksen2022integrated of varying sophistication. The former methods restrict usage to cases where all modalities are completely observed, significantly diminishing their utility in settings where this is often not the case (such as in clinical applications); the latter can lead to suboptimal performance due to the inherent limitations of imputed data. In addition, the complex and irregular temporal dynamics present in multimodal data have often been overlooked (zhang2023improving, ; tipirneni2022self, ), with existing methods often ignoring irregularity entirely soenksen2022integrated or relying on positional embedding schemes tsai2019multimodal that may not be appropriate when modalities display a varying degree of temporal irregularity. Consequently, there is a pressing need for more advanced and scalable multimodal fusion techniques that can efficiently handle a broader set of modalities, effectively manage missing and irregular data, and capture the nuanced inter-modal relationships necessary for robust and accurate prediction. We use the term FlexiModal Data to capture several of these key aspects, which haven’t been well-addressed by prior works:
@@
Table 1: We evaluated the characteristics of FuseMoE against various benchmarks. The pipeline approach soenksen2022integrated relies on a simple feature extraction scheme for each modality, followed by concatenation and classification. It doesn’t incorporate irregularities or missingness in its process, but its use of concatenation and zero-imputation for missing modalities allows it to be adapted to FlexiModal settings. Both zhang2023improving and zadeh2017tensor tackle multi-modality fusion, but as modalities increase, their method demands exponentially more cross-modal computations and significant model architecture modifications. Finally, mustafa2022multimodal presents MoE for language-image alignment, yet it also requires substantial adjustments for the more intricate and universal FlexiModal context we explore.
@@
Soenksen et al. soenksen2022integrated Data Pipeline $\geq$4
Zhang et al. zhang2023improving Modality Fusion 2
Zadeh et al. zadeh2017tensor Modality Fusion 3
Mustafa et al. mustafa2022multimodal Multimodal MoE 2
@@
FlexiModal data is most evident in clinical scenarios, where extensive monitoring results in the accumulation of comprehensive electronic health records (EHRs) for each patient. A typical EHR encompasses diverse data types, including tabular (e.g., age, demographics, gender), images (such as X-rays and magnetic resonance imaging), clinical notes, physiological time series (e.g., ECG and EEG), and vital signs (blood chemistry, heart rate, etc.). In this setting, we observe a variety of modalities, sampled with varying irregularity and a high degree of missingness and sparsity.
@@ 2 FuseMoE: Enhance Predictive Performance for FlexiModal Data
Figure 1: An example of addressing the challenge of FlexiModal Data: patients in ICUs often have extensive and irregular health status measurements over time; patients with milder conditions only require monitoring across fewer categories. FuseMoE is adept at handling inputs featuring any combination of modalities, including those with missing elements. It starts by encoding inputs using modality-specific feature extractors, followed by employing a multi-time attention mechanism (shukla2021multi, ) to address temporal irregularities. The core of FuseMoE lies the MoE Fusion Layer, where a routing mechanism is trained to categorize multimodal inputs and direct them to the appropriate combinations of MLPs. The outputs from these MLPs are weighted through a gating function, resulting in fused embeddings, which are subsequently utilized for further processing.
@@ 2.1 Sparse MoE Backbone
The main components of a sparse MoE layer are a network $G$ as a sparse gate and an expert network $E$. shazeer2017outrageously proposed a Top-$K$ gating function that takes as an input a token representation $x\in\mathbb{R}^{D}$ and then routes it to the Top-$K$ experts out of the set $\{E_{i}\}_{i=1}^{S}$. The gating network parameter $W\in\mathbb{R}^{D\times N}$ produces logits $h_{s}(x)=\mathrm{Top~{}K}(x\cdot W)$, which are normalized via Softmax:
@@
Gating Network Design The gating network’s advantage lies in its capacity to be concurrently trained with FFNs, facilitating the learning of an optimal sparse combination of experts. Essentially, by evaluating the similarity between the input token and the experts, the gating network/router optimally matches the input partition with the most suitable experts. In many cases, variations in the routing mechanism can greatly influence performance across diverse applications (li2022sparse, ). The Softmax gating is the most widely adopted across domains (riquelme2021scaling, ; shazeer2017outrageously, ). We introduce a novel Laplace gating function that offers enhanced convergence guarantees and delivers superior predictive performance, particularly in FlexiModal applications. The function is formulated as follows:
@@ 2.2 Modality and Irregularity Encoder
To encode the irregularity of sampling in each modality, we utilize a discretized multi-time attention (mTAND) module (shukla2021multi, ), which leverages a time attention mechanism (kazemi2019time2vec, ; vaswani2017attention, ) to discretize irregularly sampled observations into discrete intervals. Specifically, given a set of $l_{k}$ continuous time points, $t\in\mathbb{R}^{l_{k}}$, corresponding to the $k^{\mathrm{th}}$ dimensionality of a given modality, we employ $H$ embedding functions $\phi_{h}(\tau)$ to embed each $\tau_{k}\in t_{k}$ in a $d_{h}$ dimensional vector space (detailed definition and examples can be found in Appendix B and E). The $i^{\mathrm{th}}$ dimension of the $h^{\mathrm{th}}$ embedding is defined as
@@
where $\mathbf{Q}_{h}$ and $\mathbf{K}_{h}$ are learnable parameters. This formulation allows us to discretize univariate observations $x_{k}$ into $\gamma$ regularly-sampled bins. To model irregularity across a multivariate set of observations for a given modality with $d_{m}$ dimensions, we repeat this process for each dimension of the input. This allows us to obtain an interpolation matrix $\hat{X}_{h}=[\hat{x}_{1,h},\hat{x}_{2,h},...,\hat{x}_{d_{m},h}]\in\mathbb{R}^{\gamma\times d_{m}}$ for each of the $h$ embedding functions. We then concatenate the interpolation matrices across all $H$ embedding functions (i.e., $I=[\hat{X}_{1},\hat{X}_{2},...,\hat{X}_{h}]\in\mathbb{R}^{\gamma\times(H\cdot d_{m})}$) and employ a linear projection to achieve a final, discretized embedding for each modality, $Z\in\mathbb{R}^{\gamma\times d_{e}}$, where $d_{e}$ denotes the desired dimensionality of each modality’s representation. The discretization procedure offers a standardized approach to managing irregularly sampled time series across various input types; however, it can inevitably result in information loss. On the other hand, relying solely on the mTAND module may yield suboptimal performance due to the potentially varying sampling rates of different variables (horn2020set, ), especially in scenarios where the sample sizes are small. To mitigate this, we combine discretized outputs with continuous representations learned through the mTAND module.
Figure 2: We present three exemplary designs of the Tok-$K$ router for effective multimodal fusion, considering an input scenario with three modalities: Time-Series (TS), Text (TXT), and images (IMG). (a) The joint router design utilizes a concatenated embedding of all modalities, directing this combined input to selected experts. (b) In the modality-specific router design, each modality’s embedding is independently assigned to a shared pool of experts. (c) The third design variant also uses modality-specific routers but assigns each modality’s embedding to separate pools of experts, each pool uniquely tailored to process a specific modality type.
@@ 2.3 MoE Fusion Layer
We implement an entropy regularization loss to ensure balanced and stable expert utilization, a concept supported by various previous studies (mustafa2022multimodal, ; meister2020generalized, ; genevay2019entropy, ). It maximizes the mutual information between modalities and experts and serves as an auxiliary loss function in addition to task-specific loss. Given a total of $M$ modalities, and denoting $\mathcal{H}$ as the entropy, we define the loss function $\mathcal{E}$ as
@@
In scenarios where certain modalities are missing throughout the data trajectories, we substitute the original embedding $Z_{\mathrm{missing}}$ with a learnable embedding $\mathcal{Z}$, acting as a generic “missing indicator”. This strategy is facilitated by employing per-modality routers, which, in conjunction with entropy regularization, guide $\mathcal{Z}$ predominantly toward a specific group of less-utilized experts. The new embeddings $\mathcal{Z}$ are dynamically adjusted throughout the model training process to simultaneously minimize the task-specific loss and the entropy regularization loss. As a result, the router will assign lower weights to the experts used to process these embeddings.
@@ 3 Theoretical Contribution
Table 2: Summary of parameter estimation rates under the Softmax and Laplace gating Gaussian MoE models. The function $\widetilde{r}(\cdot)$ represents the solvability of a system of polynomial equations considered in nguyen2023demystifying while $\widetilde{r}(\cdot)\leq\bar{r}(\cdot)$ and $\widetilde{r}(2)=4$, $\widetilde{r}(3)=6$. Additionally, $\mathcal{A}^{n}_{j}:=\mathcal{A}_{j}(\widehat{G}_{n})$ denotes a Voronoi cell defined in equation 7.
@@
Softmax (nguyen2023demystifying, ) $\mathcal{O}(n^{-1/2})$ $\mathcal{O}(n^{-1/2\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/2\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$

### v4

@@ 1 Introduction
Multimodal fusion is a critical and extensively studied problem in many significant domains (Shaik et al., 2023; Yang et al., 2007; Tsai et al., 2019; Cao et al., 2023), such as sentiment analysis Han et al. (2021b); Majumder et al. (2018), image and video captioning Karpathy and Fei-Fei (2015); Johnson et al. (2016b), and medical prediction Huang et al. (2020b); Soenksen et al. (2022). Previous research has shown that embracing multimodality can improve predictive performance by capturing complementary information across modalities, outperforming single-modality approaches in similar tasks (Potamianos et al., 2003; Huang et al., 2020a). However, an ongoing challenge lies in the creation of scalable frameworks for fusing multimodal data under a variety of conditions, and in creating reliable models that consistently surpass their single-modal counterparts.
Handling a variable number of input modalities remains an open challenge in multimodal fusion, due to challenges with scalability and lack of unified approaches for addressing missing modalities. Many existing multimodal fusion methods are designed for only two modalities Han et al. (2021a); Zhou et al. (2019); Zhang et al. (2023), rely on costly pairwise comparisons between modalities Tsai et al. (2019), or employ simple concatenation approaches Soenksen et al. (2022), rendering them unable to scale to settings with a large number of input modalities or adequately capture inter-modal interactions. Similarly, existing works are either unable to handle missing modalities entirely Zhang et al. (2023); Zhan et al. (2021) or use imputation approaches Tran et al. (2017); Liu et al. (2023); Soenksen et al. (2022) of varying sophistication. The former methods restrict usage to cases where all modalities are completely observed, significantly diminishing their utility in settings where this is often not the case (such as in clinical applications); the latter can lead to suboptimal performance due to the inherent limitations of imputed data. In addition, the complex and irregular temporal dynamics present in multimodal data have often been overlooked (Zhang et al., 2023; Tipirneni and Reddy, 2022), with existing methods often ignoring irregularity entirely Soenksen et al. (2022) or relying on positional embedding schemes Tsai et al. (2019) that may not be appropriate when modalities display a varying degree of temporal irregularity. Consequently, there is a pressing need for more advanced and scalable multimodal fusion techniques that can efficiently handle a broader set of modalities, effectively manage missing and irregular data, and capture the nuanced inter-modal relationships necessary for robust and accurate prediction. We use the term FlexiModal Data to capture several of these key aspects, which haven’t been well-addressed by prior works:
@@
Table 1: We evaluated the characteristics of FuseMoE against various benchmarks. The pipeline approach (Soenksen et al., 2022) relies on a simple feature extraction scheme for each modality, followed by concatenation and classification. It doesn’t incorporate irregularities or missingness in its process, but its use of concatenation and zero-imputation for missing modalities allows it to be adapted to FlexiModal settings. Both Zhang et al. (2023) and Zadeh et al. (2017) tackle multi-modality fusion, but as modalities increase, their method demands exponentially more cross-modal computations and significant model architecture modifications. Finally, Mustafa et al. (2022) presents MoE for language-image alignment, yet it also requires substantial adjustments for the more intricate and universal FlexiModal context we explore.
@@
Soenksen et al. (2022) Data Pipeline $\geq$4
Zhang et al. (2023) Modality Fusion 2
Zadeh et al. (2017) Modality Fusion 3
Mustafa et al. (2022) Multimodal MoE 2
@@
FlexiModal data is most evident in clinical scenarios, where extensive monitoring results in the accumulation of comprehensive electronic health records (EHRs) for each patient. A typical EHR encompasses diverse data types, including tabular (e.g., age, demographics, gender), images (X-rays, magnetic resonance imaging, and photographs), clinical notes, physiological time series (ECG and EEG), and vital signs (blood chemistry, heart rate). In this setting, we observe a variety of modalities, sampled with varying irregularity and a high degree of missingness and sparsity.
@@ 2 FuseMoE: Enhance Predictive Performance for FlexiModal Data
Figure 1: An example of addressing the challenge of FlexiModal Data: patients in ICUs often have extensive and irregular health status measurements over time; patients with milder conditions only require monitoring across fewer categories. FuseMoE is adept at handling inputs featuring any combination of modalities, including those with missing elements. It starts by encoding inputs using modality-specific feature extractors, followed by employing a multi-time attention mechanism (Shukla and Marlin, 2021) to address temporal irregularities. The core of FuseMoE lies the MoE Fusion Layer, where a routing mechanism is trained to categorize multimodal inputs and direct them to the appropriate combinations of MLPs. The outputs from these MLPs are weighted through a gating function, resulting in fused embeddings, which are subsequently utilized for further processing.
@@ 2.1 Sparse MoE Backbone
The main components of a sparse MoE layer are a network $G$ as a sparse gate and an expert network $E$. Shazeer et al. (2017) proposed a Top-$K$ gating function that takes as an input a token representation $x\in\mathbb{R}^{D}$ and then routes it to the Top-$K$ experts out of the set $\{E_{i}\}_{i=1}^{S}$. The gating network parameter $W\in\mathbb{R}^{D\times N}$ produces logits $h_{s}(x)=\mathrm{Top~{}K}(x\cdot W)$, which are normalized via Softmax:
@@
Gating Network Design The gating network’s advantage lies in its capacity to be concurrently trained with FFNs, facilitating the learning of an optimal sparse combination of experts. Essentially, by evaluating the similarity between the input token and the experts, the gating network/router optimally matches the input partition with the most suitable experts. In many cases, variations in the routing mechanism can greatly influence performance across diverse applications (Li et al., 2022a). The Softmax gating is the most widely adopted across domains (Riquelme et al., 2021; Shazeer et al., 2017). We introduce a novel Laplace gating function that offers enhanced convergence guarantees and delivers superior predictive performance, particularly in FlexiModal applications. The function is formulated as follows:
@@ 2.2 Modality and Irregularity Encoder
To encode the irregularity of sampling in each modality, we utilize a discretized multi-time attention (mTAND) module (Shukla and Marlin, 2021), which leverages a time attention mechanism (Kazemi et al., 2019; Vaswani et al., 2017) to discretize irregularly sampled observations into discrete intervals. Specifically, given a set of $l_{k}$ continuous time points, $t\in\mathbb{R}^{l_{k}}$, corresponding to the $k^{\mathrm{th}}$ dimensionality of a given modality, we employ $H$ embedding functions $\phi_{h}(\tau)$ to embed each $\tau_{k}\in t_{k}$ in a $d_{h}$ dimensional vector space (detailed definition and examples can be found in Appendix B and E). The $i^{\mathrm{th}}$ dimension of the $h^{\mathrm{th}}$ embedding is defined as
@@
where $\mathbf{Q}_{h}$ and $\mathbf{K}_{h}$ are learnable parameters. This formulation allows us to discretize univariate observations $x_{k}$ into $\gamma$ regularly-sampled bins. To model irregularity across a multivariate set of observations for a given modality with $d_{m}$ dimensions, we repeat this process for each dimension of the input. This allows us to obtain an interpolation matrix $\hat{X}_{h}=[\hat{x}_{1,h},\hat{x}_{2,h},...,\hat{x}_{d_{m},h}]\in\mathbb{R}^{\gamma\times d_{m}}$ for each of the $h$ embedding functions. We then concatenate the interpolation matrices across all $H$ embedding functions (i.e., $I=[\hat{X}_{1},\hat{X}_{2},...,\hat{X}_{h}]\in\mathbb{R}^{\gamma\times(H\cdot d_{m})}$) and employ a linear projection to achieve a final, discretized embedding for each modality, $Z\in\mathbb{R}^{\gamma\times d_{e}}$, where $d_{e}$ denotes the desired dimensionality of each modality’s representation. The discretization procedure offers a standardized approach to managing irregularly sampled time series across various input types; however, it can inevitably result in information loss. On the other hand, relying solely on the mTAND module may yield suboptimal performance due to the potentially varying sampling rates of different variables (Horn et al., 2020), especially in scenarios where the sample sizes are small. To mitigate this, we combine discretized outputs with continuous representations learned through the mTAND module.
Figure 2: We present three exemplary designs of the Top-$K$ router for effective multimodal fusion, considering an input scenario with three modalities: Time-Series (TS), Text (TXT), and images (IMG). (a) The joint router design utilizes a concatenated embedding of all modalities, directing this combined input to selected experts. (b) In the modality-specific router design, each modality’s embedding is independently assigned to a shared pool of experts. (c) The third design variant also uses modality-specific routers but assigns each modality’s embedding to separate pools of experts, each pool uniquely tailored to process a specific modality type.
@@ 2.3 MoE Fusion Layer
We implement an entropy regularization loss to ensure balanced and stable expert utilization, a concept supported by various previous studies (Mustafa et al., 2022; Meister et al., 2020; Genevay, 2019). It maximizes the mutual information between modalities and experts and serves as an auxiliary loss function in addition to task-specific loss. Given a total of $M$ modalities, and denoting $\mathcal{H}$ as the entropy, we define the loss function $\mathcal{E}$ as
@@
In scenarios where certain modalities are missing throughout the data trajectories, we substitute the original embedding $Z_{\mathrm{missing}}$ with a learnable embedding $\mathcal{Z}$, acting as a generic “missing indicator”. This strategy is facilitated by employing per-modality routers, which, in conjunction with entropy regularization, guide $\mathcal{Z}$ predominantly toward a specific group of less-utilized experts. The new embeddings $\mathcal{Z}$ are dynamically adjusted throughout the model training process to minimize the task-specific loss and the entropy regularization loss. As a result, the router will assign lower weights to the experts responsible for processing these embeddings.
@@ 3 Theoretical Contribution
Problem Setup. Since the convergence analysis of MLE under the Top-K sparse gating MoE has been studied in Nguyen et al. (2024c), we will focus on examining the Laplace gating solely in the sequel. Assume that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. samples drawn from the Laplace gating Gaussian MoE of order $k_{*}$ whose conditional density function $p_{G_{*}}(Y|X)$ is
$p_{G_{*}}(Y|X)=\sum_{i=1}^{k_{*}}$ $\mathrm{softmax}(-\|W_{i}^{*}-X\|+\beta_{i}^{*})\cdot f(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*}),$ (4)
where we define for any vectors $v=(v_{i})_{i=1}^{k_{*}}$ that $\mathrm{softmax}(v_{i}):=\frac{\exp(v_{i})}{\sum_{j=1}^{k_{*}}\exp(v_{j})}$. Above, $f(\cdot|\mu,\nu)$ denotes a univariate Gaussian density function with mean $\mu$ and variance $\nu$. For ease of the presentation, we denote $G_{*}:=\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{*})\delta_{(W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})}$ as a true but unknown mixing measure associated with unknown parameters $(\beta_{i}^{*},W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})$ for $i\in\{1,2,\ldots,k_{*}\}$. In the paper, we specifically consider two settings of the true number of experts $k_{*}$: (i) Exact-specified setting: when $k_{*}$ is known; (ii) Over-specified setting: when $k_{*}$ is unknown, and we over-specify the model in equation 4 by a Laplace gating MoE model with $k>k_{*}$ experts. However, due to the space limit, we present only the latter setting, and defer the former setting to Appendix J.
Maximum Likelihood Estimation. We use the maximum likelihood method to estimate the unknown mixing measure $G_{*}$ van de Geer (2000). In particular, the MLE is given by
$\widehat{G}_{n}\in\operatorname*{arg\,max}_{G\in\mathcal{G}_{k}(\Theta)}\frac{1}{n}\sum_{i=1}^{n}\log(p_{G}(Y_{i}|X_{i})),$ (5)
where $\mathcal{G}_{k}(\Theta):=\{G=\sum_{i=1}^{k^{\prime}}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}:1\leq k^{\prime}\leq k,\ (W_{i},a_{i},b_{i},\nu_{i})\in\Theta\}$ denotes the set of all mixing measures with at most $k$ components. Given the MLE defined in equation 5, we are ready to present the main results. Before that, let us introduce some necessary notations for our analysis.
@@
Firstly, we demonstrate in Theorem 3.1 that the convergence rate of density estimation under the Laplace gating Gaussian MoE is parametric on the sample size $n$.
Theorem 3.1 (Density estimation).
The density estimation $p_{\widehat{G}_{n}}(Y|X)$ converges to the true density $p_{G_{*}}(Y|X)$ under the Total Variation distance at the following rate:
Proof of Theorem 3.1 is in Appendix K.2. The parametric rate $\mathcal{O}(\sqrt{\log(n)/n})$ of the conditional density function $p_{\widehat{G}_{n}}$ indicates that if there exists a loss function among parameters $\mathcal{D}$ such that $\mathbb{E}_{X}[V(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}(\widehat{G}_{n},G_{*})$, then we will achieve the parameter and expert estimation rates via the bound $\mathcal{D}(\widehat{G}_{n},G_{*})=\mathcal{O}(\sqrt{\log(n)/n})$.
Voronoi Loss. Following the above implication, we now define a loss function among parameters based on a notion of Voronoi cells as in Manole and Ho (2022). Given some mixing measure $G$, we distribute its components $\theta_{i}:=(W_{i},a_{i},b_{i},\nu_{i})$ to the following Voronoi cells, which are generated by the components $\theta^{*}_{j}:=(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ of the true mixing measure $G_{*}$:
for any $1\leq j\leq k_{*}$. Note that, the cardinality of the Voronoi cell $\mathcal{A}_{j}$ is exactly the number of fitted components approximating $\theta^{*}_{j}$. For ease of the presentation, let us denote $\Phi_{ij}(\rho_{1},\rho_{2},\rho_{3},\rho_{4}):=\|W_{i}-W_{j}^{*}\|^{\rho_{1}}+\|a_{i}-a_{j}^{*}\|^{\rho_{2}}+|b_{i}-b_{j}^{*}|^{\rho_{3}}+|\nu_{i}-\nu_{j}^{*}|^{\rho_{4}}$, for any $(\rho_{1},\rho_{2},\rho_{3},\rho_{4})\in\mathbb{R}^{4}$. Then, the Voronoi loss function $\mathcal{D}_{2}(G,G_{*})$ used for our analysis under the over-specified setting is given by:
$\mathcal{D}_{2}(G,G_{*}):=\sum_{j=1}^{k_{*}}\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i})-\exp(\beta^{*}_{j})\Big{|}+\sum_{j\in[k_{*}]:|\mathcal{A}_{j}|=1}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i})\Phi_{ij}(1,1,1,1)$
$\hskip 99.58464pt+\sum_{j\in[k_{*}]:|\mathcal{A}_{j}|>1}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i})\Phi_{ij}\Big{(}2,2,\bar{r}(|\mathcal{A}_{j}|),\frac{\bar{r}(|\mathcal{A}_{j}|)}{2}\Big{)}.$ (7)
The notation $\bar{r}(|\mathcal{A}_{j}|)$ stands for the minimum value of $r\in\mathbb{N}$ such that the following system of polynomial equations does not have any non-trivial solutions for the unknown variables $\{(q_{1i},q_{2i},q_{3i})\}_{i=1}^{|\mathcal{A}_{j}|}$:
$\sum_{i=1}^{|\mathcal{A}_{j}|}\sum_{\begin{subarray}{c}m_{1}+2m_{2}=s,\\ 1\leq m_{1}+m_{2}\leq r\end{subarray}}\frac{q_{3i}^{2}q_{1i}^{m_{1}}q_{2i}^{m_{2}}}{m_{1}!m_{2}!}=0,~{}\text{for each }s=1,2,\ldots,r,$ (8)
A solution to the above system is regarded as non-trivial if at least among variables $q_{1i}$ is different from zero, whereas all the variables $q_{3i}$ are non-zero. It is worth noting that the function $\bar{r}(\cdot)$ was previously studied in Ho and Nguyen (2016) to characterize the convergence behavior of parameter estimation under the location-scale Gaussian mixture models. Ho and Nguyen (2016) also gave some specific values of that function, namely $\bar{r}(2)=4$ and $\bar{r}(3)=6$. Meanwhile, they claimed that it was non-trivial to determine the value of $\bar{r}(m)$ when $m\geq 4$, and further techniques should be developed for that purpose. Since Gaussian MoE models are generalization of the Gaussian mixture models, we also involve the function $\bar{r}(\cdot)$ in our convergence analysis. Now, we provide in the following theorem the convergence rate of parameter estimation under the over-specified setting of the Laplace gating Gaussian MoE model (see also Figure 3 for the empirical convergence rates justifying the theoretical rates in Theorem 3.2).
Table 2: Parameter estimation rates under the Softmax and Laplace gating Gaussian MoE models. The function $\widetilde{r}(\cdot)$ represents the solvability of a system of polynomial equations considered in Nguyen et al. (2023) while $\widetilde{r}(\cdot)\leq\bar{r}(\cdot)$ and $\widetilde{r}(2)=4$, $\widetilde{r}(3)=6$. Additionally, $\mathcal{A}^{n}_{j}:=\mathcal{A}_{j}(\widehat{G}_{n})$ denotes a Voronoi cell defined in equation 6.
@@
Softmax (Nguyen et al., 2023) $\mathcal{O}(n^{-1/2})$ $\mathcal{O}(n^{-1/2\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/2\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$ $\mathcal{O}(n^{-1/\widetilde{r}(|\mathcal{A}^{n}_{j}|)})$